OpenStudy (anonymous):

Find the quotient. (5ab - 10b2 + 15bc) ÷ 5b a + 2b - 3c a - 2b + 3c a - 2b - 3c

OpenStudy (anonymous):

can you factor: 5ab - 10b2 + 15bc

OpenStudy (anonymous):

30b?

OpenStudy (anonymous):

LCM ( 5, 10, 15 ) = ....

OpenStudy (anonymous):

30

OpenStudy (anonymous):

20

OpenStudy (merchandize):

= (5ab - 10b² + 15bc)/5b = (5b[a - 2b + 3c])/5b = a - 2b + 3c Answer: a - 2b + 3c

OpenStudy (anonymous):

thankyou